# GWQ: Gradient-Aware Weight Quantization for Large Language Models

### v2

@@ 5 Conclusion
This paper introduces GWQ, a weight-only post-training mixed-precision quantization approach based on first-order gradients. GWQ preserves the weights corresponding to the top 1% of outliers at FP16 precision, while storing the remaining non-outlier weights in a lower-bit format. GWQ found that it is easier to obtain more accurate outliers using the first-order gradient to locate the sensitive weights of the pre-trained model than the current search for model outliers via the Hessian matrix. GWQ achieves the lowest quantization loss compared to current state-of-the-arts methods by utilizing only a single calibration sample. Considerable results are also obtained by GWQ on multimodal models, which can achieve better results on most of the zero-shot target detection tasks compared to the current stat-of-the-art methods. GWQ achieves 1.2$\times$ inference acceleration compared to the original model internship and uses less memory in the inference process.
@@ 6 Limitation
GWQ gets gradients via back propagation and ranks them accordingly. However, in models such as OPT, the activation function is ReLU, and the correct gradients cannot be computed because of gradient vanishing. Additionally, backpropagation demands substantial memory resources, a single A100 80G GPU is insufficient to handle the large computational requirements. Furthermore, since GWQ employs mixed-precision quantization, it is less hardware-friendly compared to methods like AWQ, resulting in higher inference latency after quantization. We will continue to refine and optimize the GWQ methodology.

### v3

@@ 5 Conclusion
This paper introduces GWQ, a weight-only post-training mixed-precision quantization based on first-order gradients. GWQ preserves the weights corresponding to the top 1% of outliers at FP16 bit-width while storing the remaining non-outlier weights in lower-bit. It found that it is easier to obtain more accurate outliers using the first-order gradient than Hessian matrix. GWQ achieves the lowest quantization loss compared to current SOTA methods with only one single calibration sample. In both text-only tasks and multimodel tasks, LLMs quanted by GWQ perform better than other quantization methods, which indecate that GWQ could apply to different models and tasks. Also, GWQ achieves 1.2$\times$ inference acceleration compared to the original model internship and uses less memory in the inference process.
@@ 6 Limitation
GWQ gets gradients via back propagation and ranks them accordingly. However, in models such as OPT, the activation function is ReLU, and the correct gradients cannot be computed because of gradient vanishing. Additionally, backpropagation demands substantial memory resources, and a single A100 80G GPU is insufficient to handle the large computational requirements. Furthermore, since GWQ employs mixed-precision quantization, it is less hardware-friendly compared to methods like AWQ, resulting in higher inference latency after quantization. We will continue to refine and optimize the GWQ methodology.
Appendix A Appendix
A.1 Next Token Prediction Accuracy
The accuracy evaluation of for WikiText2 and C4 datasets is based on a causal language modeling task, where given a ground truth sequence of tokens $\mathbf{y}=\{\mathbf{y}_{1},\mathbf{y}_{2},\ldots,\mathbf{y}_{T}\}$ and the predicted tokens $\hat{\mathbf{y}}=\{\hat{\mathbf{y}}_{1},\hat{\mathbf{y}}_{2},\ldots,\hat{\mathbf{y}}_{T}\}$, the accuracy is computed by shifted ground truth $\vec{\mathbf{y}}$
$\vec{\mathbf{y}}=\mathbf{y}[:,1:]$ (15)
and the most probable token of the model output logits $\mathbf{L}$, where $\mathbf{L}$ is shifted to remove the last token’s predictions
$\mathbf{L}_{\text{shifted}}=\mathbf{L}[:,:-1,:]$ (16)
and predictions $\hat{y}$ are
$\hat{\mathbf{y}}_{t}=\arg\max_{v}\mathbf{L}_{t,v}$ (17)
where $v$ is the vocabulary size. Finally, accuracy is
$\mathrm{Acc}=\frac{N_{\hat{y_{t}}=\overrightarrow{y_{t}}}}{T},$ (18)